# Stock Price Responses to Firm-Level News in Supply Chain Networks

### v4

@@ 3.1 Regression analysis
To examine the association between positive and negative sentiment in news articles about particular firms and their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window is associated with the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm $i$ listed in stock market $m$ is mentioned by news article $n$ on day $t$. Setting the time window at $w$, such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm $i$ before and after news article $n$ was disclosed on day $t$ in time window $w$, or $(\%\Delta P/P)^{\mathit{pre}}_{intw}$ and $(\%\Delta P/P)^{\mathit{post}}_{intw}$, respectively, by
@@
where $\bar{P_{i}}[s,t]$ is the average stock price of firm $i$ between day $s$ and $t$. For example, $\bar{P_{i}}[t-w,t-1]$ is the average stock price of firm $i$ for $w$ days between day t1t-1 and twt-w. Therefore, the left-hand side of equations (3) and (4) indicates the daily percentage change of the average stock price in time window $w$ in the pre- and post-news period, respectively. We take the average of stock prices over the time window to reduce their fluctuations. We skip days when stock prices are unavailable, such as Saturdays, Sundays, and holidays. Accordingly, for example, as shown in Figure 4, when news article $n$ about firm $i$ is disclosed on Friday, the 11th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average from the 8th (Tuesday) to 10th (Thursday). Similarly, the post-news change is defined as the change from the average of June 8–10 to that of June 11th, 14th (Monday), and 15th (Tuesday).
@@
Thus, our sample is restricted to firm-day observations that occur before or after news articles about the firms are disclosed. In other words, the number of observations in our sample is the cumulative number of firms mentioned in news articles during the period examined multiplied by 2. Using the pre- and post-news observations, we estimate the following estimation equation:
@@
where T is either pre or post, and $\mathit{PRE}{intw}$ and $\mathit{POST}{intw}$ are indicator variables that take a value of one if $(\%\Delta P/P)^{T}_{intw}$ represents the percentage change in the stock price of firm $i$ before or after news article $n$ is disclosed on day $t$, respectively. $\mathit{NEWS}_{int}$ represents the level of positive or negative sentiment in news article $n$ generated by FinBERT and corresponds to the sentiment indices $\text{Positive}_{int}$ and $\text{Negative}_{int}$ defined in the Methods section. $X_{mntw}$ is the percentage change of the Refinitiv index for each stock market $m$ that controls for the overall time trend of stock prices in the market, and $\mu_{s}$ is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
@@
We particularly estimate equation (5) for $w=$ 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run associations between positive news articles and stock prices. We hypothesize that a positive (negative) news article is associated with higher (lower) stock price changes in the short-run, with weaker associations in the longer run.
It should be emphasized that this analysis is comparing firms positively mentioned in news articles with those negatively mentioned, rather than comparing firms positively mentioned and those not mentioned by news articles. We focus on the sub-sample of firms mentioned by news articles and ignore those not mentioned, because the two types of firms are most likely to be intrinsically different from each other to a great extent so that comparing the two may be associated with large endogeneity biases.
@@
We further examine the association between positive or negative sentiment about firms in news articles and stock price changes of their suppliers and clients, using similar frameworks. Specifically, when we focus on upstream associations between news articles about firms and their suppliers, the variable $\mathit{NEWS}_{int}$ in equation (5) is replaced with the positive or negative sentiment indices from news articles that mention any of firm $i$’s client firms. Alternatively, when we examine the downstream associations between news articles about firms and their clients, $\mathit{NEWS}_{int}$ is replaced with the positive or negative indices of any of the suppliers of firm $i$. When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample.
We hypothesize that news sentiment about a firm is associated with stock price changes of its suppliers and clients. More precisely, news sentiment about a client firm is associated with stock price changes of its suppliers, in addition to those of the client firm itself, while news sentiment about a supplier firm is associated with stock price changes of its clients. These associations may reflect, for example, investors’ expectations about changes in demand faced by suppliers or about the performance of downstream firms, although such mechanisms are not directly tested in this study. As in the case of own-firm associations, these associations are expected to be stronger over shorter horizons and weaker over longer horizons.
@@ 4 Results
Figure 5: Coefficient estimates of the positive (panel [A]) and negative (panel [B]) sentiment indices of news articles about firms across the world on percentage changes of their own stock prices for different time windows. Panel (A) is a graphical presentation of Table 4. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the pre-news coefficient estimate for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news coefficient estimate for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news coefficients can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment but invisible because the confidence intervals are negligible compared with the point estimate.
In column (1) of Table 4, the estimated $\beta^{\mathit{pre}}_{1}$ is 0.323 and highly significant, indicating that a higher value of the positive sentiment index in a news article about a firm is associated with the percentage change in the firm’s stock price from 2 days to 1 day before the release of the news. Since the positiveness measure, $\mathit{NEWS}$ in equation (5), ranges from 0 to 1 (Figure 3), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively ($\mathit{NEWS}=1$) is 0.323% points higher than that of a firm mentioned with no positive sentiment ($\mathit{NEWS}=0$) before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive information included in news articles may be incorporated into stock prices prior to formal disclosure through non-public channels, contributing to pre-news price movements. Second, stock prices of firms positively mentioned by news articles may exhibit higher intrinsic growth than those of firms with no positive mention, leading to endogeneity biases. In the following analysis, we do not attempt to disentangle these two channels. Instead, we focus on the difference between post- and pre-news coefficient estimates, which summarizes the additional change in stock prices around public news disclosure relative to pre-existing price movements.
@@
These results in column (1) for the time window of 1 day generally hold in columns (2)-(6) for time windows of 2-5 and 30 days, while the post- and pre-news coefficient estimates and the difference between the two decrease as the time window becomes wider. By contrast, the difference between the post- and pre-news coefficient estimates is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
@@
Panels (A) and (B) of Figure 6 present coefficient estimates for positive and negative news sentiment about firms and the associated stock price changes of their suppliers. Table 5 reports the difference between post- and pre-news coefficient estimates for positive sentiment, with corresponding results for negative sentiment reported in SI Table 1. Detailed regression results are reported in SI Tables 4 and 5.
For suppliers, both pre- and post-news coefficient estimates are positive and statistically significant for positive news sentiment, with a larger post–pre difference for shorter time windows. The corresponding coefficient estimates for negative sentiment exhibit the opposite sign and similar magnitudes. Overall, the magnitude of supplier-level coefficient estimates is substantially smaller than that observed for own-firm stock prices.
@@
Figure 6: Coefficient estimates of the positive (panel (A)) and negative (panel (B) sentiment indices of news articles about firms across the world on percentage changes of stock prices of their suppliers for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news coefficients can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
@@
However, there are two notable differences between the own-firm coefficients and those for suppliers. First, the post- and pre-news coefficients for suppliers and their difference are substantially smaller than the corresponding own-firm coefficients and their difference. For example, the pre-news coefficient for suppliers for the time window of 1 day is 0.0293 (SI Table 4), while the corresponding coefficient for own firms is 0.323 (Table 4), more than 10 times larger. Similarly, the difference between the post- and pre-news coefficients, $\beta^{post}_{w}-\beta^{pre}_{w}$, is 0.033 for suppliers when the time window is one day (Table 5), while the corresponding difference for own firms is 0.600 (Table 4). Quantitatively, although positive news articles about firms are associated with stock prices of their suppliers, the corresponding coefficients for suppliers are less than 10% of those for own firms. Second, the difference between the post- and pre-news coefficients on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, indicating that the post-pre difference in coefficients for suppliers declines by 88% in 5 days. By contrast, the coefficients for own firm decline from 0.600 to 0.157 by 74%. This comparison shows that the difference between the post- and pre-news coefficients for suppliers diminishes more quickly after disclosure than the corresponding difference for firms’ own stock prices.
@@
Similarly, we examine coefficient estimates for positive and negative news sentiment in relation to stock prices of client firms and present the results in Panels (A) and (B) of Figure 7, respectively. We also show the difference between the post- and pre-news coefficient estimates for positive (Table 6) and negative (SI Table 2). Detailed regression results for clients are reported in SI Table 6 for positive sentiment and in SI Table 7 for negative sentiment. These results exhibit the same qualitative pattern as those for suppliers shown in Figure 6 and Table 5. Specifically, we find positive (negative) and significant pre- and post-news coefficients for positive (negative) news sentiment, as well as a positive (negative) and significant post–pre difference. However, the coefficients for clients are substantially smaller and decline more rapidly than those for own firms.
@@
Figure 7: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indices of news articles about firms across the world on percentage changes of stock prices of their clients for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news coefficient for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news coefficient for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news coefficients can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
@@
We further use another sample that focuses on Japanese listed firms. The results for the own-firm coefficients of positive and negative news sentiment for Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5. The detailed regression results are shown in SI Tables 8 and 9. One subtle difference is that in Figure 8 for Japan, the post-news coefficient for negative news sentiment using the time window of 2 days is larger in absolute terms than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite. This finding suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
@@
Figure 8: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indices of news articles about Japanese firms on percentage changes of their own stock prices for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news effects can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
In addition, we examine coefficients for news sentiment about Japanese firms on stock prices of their suppliers and clients. Detailed regression results for suppliers are reported in SI Tables 10 and 11, and those for clients are reported in SI Tables 12 and 13.
A notable difference between the Japanese supply chain data and the global supply chain data is that the former provides dense coverage of domestic supply chain links based on firm-level surveys, whereas the latter identifies major customer–supplier links primarily from financial statements (see Data Section). Consistent with this difference in coverage, the average number of supply chain links per listed firm in 2016 is 6.4 in Japan and 3.3 in the global sample (Tables 1 and 2).
@@
Second, the post-news coefficient for positive news sentiment is smaller than its pre-news coefficient (Panel (A) of Figure 9), although the post-news coefficient is always larger in size than the pre-news coefficient in the global sample. This finding implies that positive information about a firm may be incorporated into prices prior to disclosure, for example suggestions of financial companies, SNS, and word of mouth, and that such information is associated with higher stock prices of the firm’s suppliers. However, once the positive information is disclosed in a news article, the magnitude of the coefficients for suppliers is smaller than that in the pre-news period. In other words, for Japanese firms, the difference between the post- and pre-news effects on suppliers is negative, indicating that the pre-news effect exceeds the post-news effect.
@@
Within this common pattern, however, there are differences across time windows. In particular, for Japanese firms, the pre-news effect of negative news for the one-day time window is smaller in absolute value than that for longer windows of two to five days, whereas the global results show an increase in the absolute value of the pre-news effect as the window size decreases.
@@
Figure 9: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indices of news articles about Japanese firms on percentage changes of their suppliers’ stock prices for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news effects can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
We further examine coefficients for news sentiment on client firms and present the results in Figure 10. Detailed regression results are reported in SI Tables 12 and 13. Compared with the corresponding results for firms across the world shown in Figure 7, we find that the same three differences observed for suppliers also apply to clients. First, the average magnitude of the effects is generally larger for Japanese firms than for firms in the global sample. Second, for positive news, the difference between the post- and pre-news effects on clients is negative in Japan. Third, for negative news, the pre-news effect for the one-day time window is smaller than that for longer time windows.
@@
Figure 10: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indices of news articles about Japanese firms on percentage changes of their clients’ stock prices for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news effects can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.

### v5

@@ 3.1 Regression analysis
To examine the association between positive and negative sentiment in news articles about particular firms and their own stock prices, we apply the positive and negative sentiment indexes constructed above in regression analysis. Specifically, we focus on firms mentioned by news articles and examine whether the stock price change rate in a time window is associated with the level of positiveness or negativeness of the article in the pre- and post-disclosure periods.
Suppose that firm $i$ listed in stock market $m$ is mentioned by news article $n$ on day $t$. Setting the time window at $w$, such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm $i$ before and after news article $n$ was disclosed on day $t$ in time window $w$, or $(\%\Delta P/P)^{\mathit{pre}}_{intw}$ and $(\%\Delta P/P)^{\mathit{post}}_{intw}$, respectively, by
@@
where $\bar{P_{i}}[t_{1},t_{2}]$ is the average stock price of firm $i$ between day $t_{1}$ and $t_{2}$. For example, $\bar{P_{i}}[t-w,t-1]$ is the average stock price of firm $i$ for $w$ days between day twt-w and t1t-1. Therefore, the left-hand side of equations (3) and (4) indicates the daily percentage change of the average stock price in time window $w$ in the pre- and post-news period, respectively. We take the average of stock prices over the time window to reduce their fluctuations. We skip days when stock prices are unavailable, such as Saturdays, Sundays, and holidays. Accordingly, for example, as shown in Figure 4, when news article $n$ about firm $i$ is disclosed on Friday, the 11th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average from the 8th (Tuesday) to 10th (Thursday). Similarly, the post-news change is defined as the change from the average of June 8–10 to that of June 11th, 14th (Monday), and 15th (Tuesday).
@@
Accordingly, the unit of observation in our analysis is a firm–news-article pair. For each firm mentioned in a news article, we construct two observations corresponding to the pre- and post-disclosure periods. As a result, the total number of observations equals twice the cumulative number of firm mentions in news articles over the sample period.
Using the pre- and post-news observations, we estimate the following estimation equation:
@@
where T is either pre or post, and $\mathit{PRE}{intw}$ and $\mathit{POST}{intw}$ are indicator variables that take a value of one if $(\%\Delta P/P)^{T}_{intw}$ represents the percentage change in the stock price of firm $i$ over the length of the time window $w$ before or after news article $n$ is disclosed on day $t$, respectively. $\mathit{NEWS}_{int}$ represents the level of positive or negative sentiment in news article $n$ generated by FinBERT and corresponds to the sentiment indexes $\text{Positive}_{int}$ and $\text{Negative}_{int}$ defined in the Methods section. $X_{mntw}$ is the percentage change of the Refinitiv index for each stock market $m$ that controls for the overall time trend of stock prices in the market, and $\mu_{s}$ is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
@@
We particularly estimate equation (5) for $w=$ 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run associations between positive news articles and stock prices. We hypothesize that a positive (negative) news article is associated with higher (lower) stock price changes in the short-run, with weaker associations in the longer run. The analysis is conducted on the sub-sample of firms mentioned in news articles, as firms that are mentioned and those that are not mentioned are likely to differ systematically in ways that are difficult to control for.
@@
We further examine the association between positive or negative sentiment about firms in news articles and stock price changes of their suppliers and clients, using analogous frameworks. Specifically, when we focus on upstream associations between news articles about firms and their suppliers, the variable $\mathit{NEWS}_{int}$ in equation (5) is replaced with the positive or negative sentiment indixes from news articles that mention any of firm $i$’s client firms. Alternatively, when we examine the downstream associations between news articles about firms and their clients, $\mathit{NEWS}_{int}$ is replaced with the positive or negative indixes of any of the suppliers of firm $i$. When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample.
We hypothesize that news sentiment about a firm is associated with stock price changes of its suppliers and clients. These associations may reflect, for example, investors’ expectations about changes in demand faced by suppliers or about the performance of downstream firms, although such mechanisms are not directly tested in this study. As in the case of own-firm associations, these associations are expected to be stronger over shorter horizons and weaker over longer horizons.
@@ 4 Results
Figure 5: Coefficient estimates of the positive (panel [A]) and negative (panel [B]) sentiment indexes of news articles about firms across the world on percentage changes of their own stock prices for different time windows. Panel (A) is a graphical presentation of Table 4. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the pre-news coefficient estimate for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news coefficient estimate for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news coefficients can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment but invisible because the confidence intervals are negligible compared with the point estimate.
In column (1) of Table 4, the estimated $\beta^{\mathit{pre}}_{1}$ is 0.323 and highly significant, indicating that a higher value of the positive sentiment index in a news article about a firm is associated with the percentage change in the firm’s stock price from two days before to one day before the release of the news. Since the positiveness measure, $\mathit{NEWS}$ in equation (5), ranges from 0 to 1, the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively ($\mathit{NEWS}=1$) is 0.323% points higher than that of a firm mentioned with no positive sentiment ($\mathit{NEWS}=0$) before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive information, i.e., actual positive incident is incorporated into stock prices prior to disclosure of news. Second, stock prices of firms positively mentioned by news articles exhibits higher intrinsic growth than those of firms with no positive mention, leading to endogeneity biases. In the following analysis, we do not attempt to disentangle these two channels. Instead, we focus on the difference between post- and pre-news coefficient estimates, which summarizes the additional change in stock prices around public news disclosure relative to pre-existing price movements.
@@
These results in column (1) for the time window of 1 day generally hold in columns (2)–(6) for time windows of 2–5 and 30 days, while the post- and pre-news coefficient estimates and the difference between the two decrease as the time window becomes wider. By contrast, the difference between the post- and pre-news coefficient estimates is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
@@
Panels (A) and (B) of Figure 6 present coefficient estimates for positive and negative news sentiment about firms and the associated stock price changes of their suppliers. Detailed regression results are reported in SI Table 4 for positive sentiment and SI Table 5 for negative sentiment, respectively. Table 5 reports the difference between post- and pre-news coefficient estimates for positive sentiment, with corresponding results for negative sentiment reported in SI Table 1.
For suppliers, both pre- and post-news coefficient estimates are positive (except for the -365-day window) and statistically significant for positive news sentiment, with a larger post–pre difference for shorter time windows. The corresponding coefficient estimates for negative sentiment exhibit the opposite sign and similar magnitudes.
@@
Figure 6: Coefficient estimates of the positive (panel (A)) and negative (panel (B) sentiment indexes of news articles about firms across the world on percentage changes of stock prices of their suppliers for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news coefficients can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
@@
There are two notable differences between the own-firm coefficients and those for suppliers. First, the post- and pre-news coefficients for suppliers and their difference are substantially smaller than the corresponding own-firm coefficients and their difference. For example, the pre-news coefficient for suppliers for the time window of 1 day is 0.0293 (SI Table 4), while the corresponding coefficient for own firms is 0.323 (Table 4), more than 10 times larger. Similarly, the difference between the post- and pre-news coefficients, $\beta^{post}_{w}-\beta^{pre}_{w}$, is 0.033 for suppliers when the time window is one day (Table 5), while the corresponding difference for own firms is 0.600 (Table 4). Quantitatively, although positive news articles about firms are associated with stock prices of their suppliers, the corresponding coefficients for suppliers are less than 10% of those for own firms. Second, the difference between the post- and pre-news coefficients on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, indicating that the post-pre difference in coefficients for suppliers declines by 88% in 5 days. By contrast, the coefficients for own firm decline from 0.600 to 0.157 by 74%. This comparison shows that the difference between the post- and pre-news coefficients for suppliers diminishes more quickly after disclosure than the corresponding difference for firms’ own stock prices.
@@
Similarly, we examine coefficient estimates for positive and negative news sentiment in relation to stock prices of client firms and present the results in Panels (A) and (B) of Figure 7, respectively. Detailed regression results for clients are reported in SI Table 6 for positive sentiment and in SI Table 7 for negative sentiment. We also show the difference between the post- and pre-news coefficient estimates for positive (Table 6) and negative (SI Table 2). These results exhibit the same qualitative pattern as those for suppliers shown in Figure 6 and Table 5. Specifically, we find positive (negative) and significant pre- and post-news coefficients for positive (negative) news sentiment, as well as a positive (negative) and significant post–pre difference. However, the coefficients for clients are substantially smaller and decline more rapidly than those for own firms.
@@
Figure 7: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indexes of news articles about firms across the world on percentage changes of stock prices of their clients for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news coefficient for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news coefficient for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news coefficients can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
@@
We further use another sample that focuses on Japanese listed firms. The results for the own-firm coefficients of positive and negative news sentiment for Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5. The detailed regression results are shown in SI Tables 8 for positive and 9 for negative, respectively. One subtle difference is that in Figure 8 for Japan, the post-news coefficient for negative news sentiment using the time window of 2 days is larger in absolute terms than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite. This result suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
@@
Figure 8: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indexes of news articles about Japanese firms on percentage changes of their own stock prices for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news effects can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
In addition, we examine coefficients for news sentiment about Japanese firms on stock prices of their suppliers and clients. Detailed regression results for suppliers are reported in SI Tables 10 and 11, and those for clients are reported in SI Tables 12 for positive and 13 for negative, respectively.
@@
Second, the post-news coefficient for positive news sentiment is smaller than its pre-news coefficient (Panel (A) of Figure 9), although the post-news coefficient is always larger in size than the pre-news coefficient in the global sample. This result implies that positive information about a firm is incorporated into prices prior to disclosure. However, once the positive information is disclosed in a news article, the magnitude of the coefficients for suppliers is smaller than that in the pre-news period. In other words, for Japanese firms, the difference between the post- and pre-news effects on suppliers is negative, indicating that the pre-news effect exceeds the post-news effect.
@@
Within this common pattern in negative cases between the entire world and Japan, however, there are differences across time windows. In particular, for Japanese firms, the pre-news effect of negative news for the one-day time window is smaller in absolute value than that for longer windows of two to five days, whereas the global results show an increase in the absolute value of the pre-news effect as the window size decreases.
@@
Figure 9: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indexes of news articles about Japanese firms on percentage changes of their suppliers’ stock prices for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news effects can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.
We further examine coefficients for news sentiment on client firms and present the results in Figure 10. Detailed regression results are reported in SI Tables 12 for positive and 13 for negative, respectively. Compared with the corresponding results for firms across the world shown in Figure 7, we find that the same three differences observed for suppliers also apply to clients. First, the average magnitude of the effects is generally larger for Japanese firms than for firms in the global sample. Second, for positive news, the difference between the post- and pre-news effects on clients is negative in Japan. Third, for negative news, the pre-news effect for the one-day time window is smaller than that for longer time windows.
@@
Figure 10: Coefficient estimates for the positive (panel (A)) and negative (panel (B)) sentiment indexes of news articles about Japanese firms on percentage changes of their clients’ stock prices for different time windows. When the value of the horizontal axis is negative and w-w, the dot above the value indicates the point estimate of the pre-news effect for time window $w$ ($\beta^{\mathit{pre}}_{w}$). When it is positive and $w$, the dot indicates the post-news effect for time window $w$ ($\beta^{\mathit{post}}_{w}$). The color of the dots for time windows $w$ and w-w is set to be the same so that the post- and pre-news effects can be easily compared. The confidence interval at the 5% level associated with each point estimate (dot) is shown by a vertical segment.